After the affordable care act was passed into law, states debated on whether they should adopt the Medicaid expansion. Was a state’s decision to adopt the Medicaid expansion influenced by its majority party affiliation? To test this claim, the following data is available.

What criticism might you have of the data analysis? Researchers might take issue with the variables chosen, or the setup of this particular test. There are no wrong answers, here, but your reasoning should be supported by statistical theory or supporting evidence.

States

Medicaid Expanded

Medicaid Not Expanded

Voted Democrat in 2012 presidential election

22

4

Voted Republican in 2012 presidential election

6

18
Answer #1
Given table data is as below
MATRIX col1 col2 TOTALS
row 1 22 4 26
row 2 6 18 24
TOTALS 28 22 N = 50
------------------------------------------------------------------

calculation formula for E table matrix
E-TABLE col1 col2
row 1 row1*col1/N row1*col2/N
row 2 row2*col1/N row2*col2/N
------------------------------------------------------------------

expected frequencies calculated by applying E - table matrix formula
E-TABLE col1 col2
row 1 14.56 11.44
row 2 13.44 10.56
------------------------------------------------------------------

calculate chisquare test statistic using given observed frequencies, calculated expected frequencies from above
Oi Ei Oi-Ei (Oi-Ei)^2 (Oi-Ei)^2/Ei
22 14.56 7.44 55.354 3.8018
4 11.44 -7.44 55.354 4.8386
6 13.44 -7.44 55.354 4.1186
18 10.56 7.44 55.354 5.2418
ᴪ^2 o = 18.0008
------------------------------------------------------------------

set up null vs alternative as
null, Ho: no relation b/w X and Y OR X and Y are independent
alternative, H1: exists a relation b/w X and Y OR X and Y are dependent
level of significance, α = 0.05
from standard normal table, chi square value at right tailed, ᴪ^2 α/2 =3.8415
since our test is right tailed,reject Ho when ᴪ^2 o > 3.8415
we use test statistic ᴪ^2 o = Σ(Oi-Ei)^2/Ei
from the table , ᴪ^2 o = 18.0008
critical value
the value of |ᴪ^2 α| at los 0.05 with d.f (r-1)(c-1)= ( 2 -1 ) * ( 2 - 1 ) = 1 * 1 = 1 is 3.8415
we got | ᴪ^2| =18.0008 & | ᴪ^2 α | =3.8415
make decision
hence value of | ᴪ^2 o | > | ᴪ^2 α| and here we reject Ho
ᴪ^2 p_value =0


ANSWERS
---------------
null, Ho: no relation b/w X and Y OR X and Y are independent
alternative, H1: exists a relation b/w X and Y OR X and Y are dependent
test statistic: 18.0008
critical value: 3.8415
p-value:0
decision: reject Ho

we have enough evidence to support the claim that state’s decision to adopt the Medicaid expansion influenced by its majority party affiliation.